At one point in a pipeline, the water’s speed is 3.50 [m/s] and the gauge pressure is 5.00 × 10^4[Pa]. Find the gauge pressure at a second point in the line, 13.0 [m] lower than the first, if the pipe diameter at the second point is twice that at the first.

Note: STEP BY STEP DON"T JUST GIVE THE ANSWER!!!


1
Expert's answer
2021-02-01T15:03:15-0500
"A_1v_1=A_2v_2\to d_1^2v_1=d_2^2v_2"

"p_1+0.5\rho v_1^2+\rho gh=p_2+0.5\rho v_2^2\\
p_1+0.5\rho v_1^2+\rho gh=p_2+0.5\rho v_1^2\frac{d_1^4}{d_2^4}\\\\
5\cdot10^4+0.5(1000) 3.5^2+(1000)(9.8)(13)\\=p_2+0.5(1000)3.5^2\frac{1}{2^4}\\p_2=1.83\cdot10^5\ Pa"